OpenStudy (anonymous):

Find the minimal polynomial of the matrix and determine whether it is diagonalizable.

OpenStudy (anonymous):

\[\left[\begin{matrix}1 & 3 & 4 \\ 2 & 2 & 4 \\ 0 & 0 & -1\end{matrix}\right]\]

OpenStudy (anonymous):

I have found the minimal polynomial which is \[m(x)=(x-4)(x+1)\]So how do I tell if it is diagonalizable or not?

OpenStudy (anonymous):

I don't know, but try reading up on the literature in Linear Algebra to find out! :-D